# MetaTrading: An Immersion-Aware Model Trading Framework for Vehicular Metaverse Services

### v2

@@ III System Model
We first present the system overview of the immersion-aware model trading framework in Section III-A. Next, we introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the learning model in Section III-C.
@@ III-A System Overview
Fig. 5 illustrates the workflow of our immersion-aware model trading framework, designed to enhance the immersive and interactive AR experiences of MUs. MUs can contribute learning models to the AR services of MSPs for better immersive experiences. For example, services such as object detection can be improved by training on images of streets, pedestrians, and so on, captured by vehicles’ cameras. In this way, AR services can be better applied to visible driving, with more accurate and timely hazard warnings from windscreens. The framework mainly consists of MSPs owned by different companies (e.g., Meta, Nissan, and Google), an infrastructure layer, and an interaction layer with the MUs that enjoy vehicular metaverse services. MSPs are powered by various technologies supporting vehicular metaverse services, such as AI engines, AR/VR engines, metaverse trading engines. The infrastructure layer with base stations (equipped with caches and edge servers) provides the basis of 5G/6G communication services for the interactions among MUs and MSPs. MUs play essential roles as both consumers and contributors to vehicular metaverse services.
TABLE I: Definitions of Notation
@@
The framework is a multi-engine configuration, and this paper focuses on two components: the trading engine and the AI engine of the MSP. The trading engine is responsible for incentivizing MUs to participate in mutually beneficial collaborations with MSPs, providing them with a trading platform to make their decisions (i.e., computational and communication resource allocation of MUs and reward decisions of MSPs). Then, each MU deciding to contribute utilizes the FL mechanisms integrated into AI engines to provide learning models for MSPs under the strategic guidance of the trading results. The specific workflow of the framework involves the following two phases.
Phase I (Incentive process): Initially, MSPs’ AI engines broadcast their global models to all MUs in the vehicular metaverse [^6] The FL global models owned by different MSPs we consider could be object detection or classification models for intelligent terrain mapping, intuitive road safety, visible driving, etc. Note that it is feasible to train on multiple tasks with the same local data. For instance, for the same image, intuitive road safety focuses on pedestrians and vehicles, while intelligent terrain mapping focuses on landmarks and routes. (① in Fig. 5). Then, the trading engines of MSPs determine their digital currency prices per IoM, i.e., rewards to MUs, and broadcast them to MUs (②). Next, MUs determine their allocation of computational and communication resources for contributing learning models to various MSPs based on the IoM, costs, and rewards given by MSPs (③). Based on the responses from MUs, the trading engines of MSPs adjust the digital currency prices to maximize their own utilities. Steps ② and ③ are repeated until an agreement is reached among MUs and MSPs.
Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computational resources to local training based on different global models from MSPs and generates local learning models for different MSPs (④). Then, MUs upload their updated learning models to the corresponding AI engines of MSPs by consuming communication resources (⑤). Following this, AI engines aggregate models from all MUs to update their global models and transmit them to their AR/VR engines for utilization (⑥). Finally, the AR/VR engines provide enhanced AR services to MUs (⑦). Steps ④-⑦ are repeated until the guidance time $T$ ends. To ease the presentation, we summarize some important notations in Table I.
@@ III-B FL Framework Mechanism Adopted
Here, we consider a set of =$\mathcal{M}=\left\{1,\ldots,M\right\}$ MUs with the capacities of local computation and a set of $\mathcal{N}=\left\{1,\ldots,N\right\}$ MSPs with FL synchronous tasks [^7] That is, the update time of the global model is limited by the slowest MU, i.e., the AI engine of the MSP must wait to receive updated models from all MUs before performing model aggregation. The advantages of synchronous FL are high model accuracy and fast convergence [29].. Each MSP initiates an FL synchronous task with a virtual deadline $\tau_{n}(n\in\mathcal{N})$. The MUs can join FL to generate local learning models and contribute to enhancing vehicular metaverse services. As shown in Fig. 4, MU $m$ provides learning models to different MSPs simultaneously for a given time period $T$, meaning multiple FL tasks can be processed in parallel. Moreover, the data collected at the $(r-1)$-th round is used as the local data for training at the $r$-th round. Without loss of generality, we consider that the trading decisions among MUs and MSPs are made at the beginning of the period $T$.
@@
In detail, let $X_{mn}$ be the set of data samples for task $n$ of MU $m$ as a set of input-output pairs $\sum_{i=1}^{|X_{mn}|}(y_{i},z_{i})$, where $y_{i}$ is an input sample vector with $d$ features, $z_{i}$ is the labeled output (i.e., ground truth) value of sample $y_{i}$, and $|X_{mn}|$ is the size of set $X_{mn}$. The data, such as images of streets and pedestrians, can be generated by high-definition cameras inside and outside the vehicles. The steps involved in each iteration $r\in\left\{1,2,\ldots,R\right\}$ are as follows:
@@
  • Uploading and Aggregation: MUs transmit their updated local model $\jmath_{mn}^{(r)}$ to the AI engine of MSP $n$. After receiving all the updated models, the MSP $n$’s AI engine aggregates them through the following weighted average:
@@
III-C Immersion Metric of Learning Model
MSPs determine the rewards for MUs based on the immersion of the learning model (IoM) metric. Correspondingly, MUs need to decide the IoMs of the models provided to MSPs by allocating their computational and communication resources to maximize the benefits. This process highlights the key role of “IoM” in the interactions between the two parties, serving as both an evaluation criterion and a basis for decision-making. Unlike traditional FL and incentive mechanisms, in vehicular metaverse scenarios with immersion requirements, both the contribution and freshness of the learning model need to be considered [^8] For example, an MSP may receive fresher learning models but with lower contributions, or less fresh learning models but with higher contributions. , which will collectively affect the immersive experience of AR services, i.e., whether virtual objects can be placed in the physical world accurately and promptly. With the above consideration, we propose the IoM, a metric that jointly considers the contribution prediction $I_{mn}$ and the age of information (AoI) $\Delta_{mn}$ [30] to measure the value of a learning model that MU $m$ brought to the vehicular metaverse service of MSP $n$, denoted as
@@ III-C Immersion Metric of Learning Model
III-C1 Contribution Prediction $I$
The contribution prediction $I_{mn}$ from MU $m$ to MSP $n$ is determined by the accuracy $\theta_{m}$ of the learning model, the total amount of training data $\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|$, and the potential value $\omega_{mn}$ of local data. Prior work [31] characterized the contribution of a learning model in terms of a logarithmic function of the amount of local training data. However, evaluating the contribution solely by the training data size is one-sided, as there may be a large amount of redundant data or the validity of the learning model cannot be guaranteed. It is more practical to comprehensively evaluate the amount of training data, model accuracy, and the potential value of local data; thus, we denote $I_{mn}$ as
@@ III-C1 Contribution Prediction $I$
where $\lfloor\frac{T}{\tau_{n}}\rfloor$ is the number of iterations that the task of MSP $n$ can be performed within $T$. $\epsilon$ and $\eta$ are system parameters from experiments as obtained in [32]. $\theta_{m}\in\left(0,1\right)$ characterizes the accuracy of local training as decided by the MU. Here, $\theta\rightarrow 0$ means that a high-accuracy learning model is available, while $\theta\rightarrow 1$ implies that the accuracy of the learning model is low. The local data potential value $\omega_{mn}$ describes the difference between model predictions and true labels, defined as
@@
where $\hat{z}_{i}$ denotes the inferred label from the current global model and $z_{i}$ is its ground-truth label. If the value of $\omega_{mn}$ is large, it indicates that the current global model is not working well on local data or there are unseen samples. Therefore, updating the model by training on these data may lead to better performance of the global model. If the value of $\omega_{mn}$ is small, it means that the current data has little potential to further improve the global model since the knowledge of the data has already been learned or the data size is small. Note that to shorten the decision time, we use the potential value of the initial local data to be a proxy for the average potential value of all data within $T$. This is a reasonable approach when $T$ is not too long.
Figure 5: The AoI evolution of the models sent from MU $m$ to MSP $n$.
@@ III-C2 Age of Information $\Delta$
The whole FL process involves model training, uploading, and waiting, which together constitute a complete cycle of model updating. From the perspective of MSP $n$, the age evolution of the model contributed by MU $m$ is shown in Fig. 5. Here, the filled circles denote the time instances when MU $m$ starts training the learning model. The intersections show the time instances when MSP $n$ receives the corresponding updated model, at which point the AoI drops to the lowest. We denote that at time $t$, the generation time of the last update received by MSP $n$ from MU $m$ is $s(t)$. Then, the instantaneous AoI of MU $m$’s learning model measured at MSP $n$, $\Delta_{mn}(t)$, is
@@
which shows the elapsed time since the generation of the latest updated model. A smaller instantaneous AoI means that the latest model received is fresher at that time. To obtain a comprehensive understanding of the overall freshness of model updates over a certain period, we use average AoI to evaluate the freshness of the models. With the aid of Figure 5, the time-averaged AoI of the learning model from MU $m$ to MSP $n$ can be expressed as
[…]
@@
where ${Q_{mn}}^{(r)}$ is the $r$-th trapezium under the curve. A lower average AoI indicates that the updated models are generally fresher over a long period.
Since MU $m$ constantly collects data throughout the FL process, the amount of training data collected for training in each round is $|X_{mn}|=x_{m}\tau_{n}$, where $x_{m}$ denotes the data (number of floats) collected per unit time. The cumulative time $T_{mn}^{c}$ for local training is determined by the amount of training data, computational resource $f_{mn}$ allocated to MSP $n$, and its local accuracy threshold $\theta_{m}$, which is equivalent to
@@
Furthermore, the upload time $T_{mn}^{t}$ is directly related to the number of model parameters $b_{mn}$ and the resource $B_{mn}$ (i.e., uplink bandwidth) used to communicate between MU $m$ and MSP $n$, defined as
@@
where $\varsigma_{mn}=\frac{p^{t}_{m}g_{mn}}{\sigma^{2}}$ denotes the signal-to-noise ratio (SNR) [34] for the communication between MU $m$ and MSP $n$. $g_{mn}$ and $p^{t}_{m}$ are the corresponding channel gain and the transmission power of MU $m$, respectively. $\sigma^{2}$ is additive white Gaussian noise (AWGN).
@@ IV-A Utility of MU
$\max\,\,\Phi_{m}(\bm{f}_{m},\bm{B}_{m})=\sum_{n=1}^{N}(p_{mn}V_{mn}-C_{mn}),$ (11)
$s.t.$ $C1:\sum_{n=1}^{N}f_{mn}<f_{m}^{max},\,\,\sum_{n=1}^{N}B_{mn}<B_{m}^{max},$
$C2:\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}\leq\tau_{n},$
@@
where $S_{m}$ indicates the minimum computational resource required to enjoy other basic services. Constraints C1 and C3 indicate that the MU’s total computational and communication resources are limited and cannot be fully used to contribute learning models, given the MU’s demand for basic services. Constraint C2 ensures that the total time for local training and uploading cannot exceed the time constraint $\tau_{n}$ for each round. The goal of each MU is to choose the optimal allocation of computational $\bm{f}_{m}$ and communication $\bm{B}_{m}$ resources to maximize its utility. Note that $V_{mn}$ must be greater than 00. This is because if $V_{mn}\leq 0$ and MU $m$ is involved in contributing learning models, the utility is calculated as $p_{mn}V_{mn}-C_{mn}\leq 0$, resulting in MU $m$ choosing not to participate in the trading.
@@ IV-B Utility of MSP
Each MSP has a gain function $\psi$ associated with the IoM to measure the benefits of all the learning models from MUs to the service. In this paper, we adopt $\psi_{n}=\mu_{n}\ln(1+\sum_{m=1}^{M}V_{mn})$ as a monotonically increasing, differentiable, strictly concave IoM function, which is a simplification of the widely adopted function [35]. Let $\bm{p}_{n}\triangleq\left(p_{1n},\ldots,p_{mn}\right)$ and $\bm{p}_{\textrm{-}n}$ represent the reward profile of MSP $n$ and all other MSPs except MSP $n$, respectively. Then, the optimization problem for the MSP is defined as
@@
$s.t.\,\,p_{mn}>0,n\in\mathcal{N},m\in\mathcal{M},$
where $\mu_{n}$ refers to the profit conversion coefficient [36] from IoM, which is adjusted according to the AR services offered by the different MSPs. Note that the other MSPs’ decisions $\bm{p}_{\textrm{-}n}$ are captured by the MUs’ responses. Moreover, when MSP $n$ determines its reward $\bm{p}_{n}$ for different MUs, the MSP needs to consider the rewards offered by other MSPs (i.e., $\bm{p}_{\textrm{-}n}$ ) as well as the strategies of all MUs (i.e., $(\bm{f}_{m},\bm{B}_{m}),m\in\mathcal{M}$). This thereby leads to reward competition among MSPs. Therefore, the optimization among MSPs can be considered as a non-cooperative game, termed a multi-MSP rewarding game $\Omega$ as follows.

### v3

@@ III System Model
We first present the system overview of the immersion-aware model trading framework in Section III-A. Next, we introduce the FL mechanism adopted by our framework in Section III-B and design the immersion metric of the local model in Section III-C.
@@ III-A System Overview
Fig. 5 illustrates the workflow of our immersion-aware model trading framework, designed to enhance the immersive and interactive AR experiences of MUs. MUs can contribute local models to the AR services of MSPs for better immersive experiences. For example, services such as object detection can be improved by training on street and pedestrian images captured by vehicles’ cameras. In this way, AR services can be applied more effectively to visible driving, with more accurate and timely hazard warnings from windscreens. The framework mainly consists of MSPs owned by different companies (e.g., Meta, Nissan, and Google), an infrastructure layer, and an interaction layer with the MUs that enjoy vehicular metaverse services. MSPs are powered by various technologies supporting vehicular metaverse services, such as AI engines, AR/VR engines, and metaverse trading engines. The infrastructure layer with base stations (equipped with caches and edge servers) provides the basis of 5G/6G communication services for the interactions among MUs and MSPs. MUs play essential roles as both consumers and contributors to vehicular metaverse services.
TABLE II: Definitions of Notation
@@
The framework adopts a multi-engine architecture, and this paper focuses on two components: the trading engine and the AI engine of the MSP. The trading engine is responsible for incentivizing MUs to participate in mutually beneficial collaborations with MSPs, providing them with a trading platform to make their decisions (i.e., computational and communication resource allocation of MUs and reward decisions of MSPs). Then, each MU deciding to contribute utilizes the FL mechanisms integrated into AI engines to provide local models for MSPs under the strategic guidance of the trading results. The specific workflow of the framework involves the following two phases.
Phase I (Incentive process): Initially, MSPs’ AI engines broadcast their global models to all MUs in the vehicular metaverse [^6] The FL global models owned by different MSPs we consider could be object detection or classification models for intelligent terrain mapping, intuitive road safety, visible driving, etc. Note that it is feasible to train on multiple tasks with the same local data. For instance, for the same image, intuitive road safety focuses on pedestrians and vehicles, while intelligent terrain mapping focuses on landmarks and routes. (① in Fig. 5). Then, the trading engines of MSPs determine their digital currency prices per IoM, i.e., rewards to MUs, and broadcast them to MUs (②). Next, each MU determines its allocation of computational and communication resources for contributing local models to various MSPs based on the IoM, costs, and rewards given by MSPs (③). Based on the responses from MUs, the trading engines of MSPs adjust the digital currency prices to maximize their own utilities. Steps ② and ③ iterate until an agreement is reached between MUs and MSPs.
Phase II (FL process): Each MU, guided by decisions obtained through tradings, allocates computational resources to local training based on different global models from MSPs and generates local models for different MSPs (④). Then, MUs upload their local updates to the corresponding AI engines of MSPs by consuming communication resources (⑤). Following this, AI engines aggregate local updates from all MUs to update their global models and transmit them to their AR/VR engines for utilization (⑥). Finally, the AR/VR engines provide enhanced AR services to MUs (⑦). Steps ④-⑦ are repeated until the guidance time $T$ ends. To ease the presentation, we summarize some important notations in Table II.
@@ III-B FL Framework Mechanism Adopted
We consider a set of $\mathcal{M}=\left\{1,\ldots,M\right\}$ MUs equipped with local computational capabilities and a set of $\mathcal{N}=\left\{1,\ldots,N\right\}$ MSPs with FL synchronous tasks [^7] That is, the update time of the global model is limited by the slowest MU, i.e., the AI engine of the MSP must wait to receive local updates from all MUs before performing model aggregation. The advantages of synchronous FL are high model accuracy and fast convergence [30].. Each MSP initiates an FL synchronous task with a virtual deadline $\tau_{n}\ (n\in\mathcal{N})$. The MUs can participate in FL to generate local models and contribute to enhancing vehicular metaverse services. As shown in Fig. 4, MU $m$ provides local models to different MSPs simultaneously for a given time period $T$, meaning multiple FL tasks can be processed in parallel. Moreover, the data collected at the (r1)(r-1)-th round are used for local training at the $r$-th round.
Without loss of generality, we consider that the trading decisions among MUs and MSPs are made at the beginning of period $T$.
@@
Let $X_{mn}=\{(y_{i},z_{i})\}_{i=1}^{|X_{mn}|}$ be the set of input-output pairs sampled from MU $m$ for task $n$, where $y_{i}$ is the input with $d$ features and $z_{i}$ is its corresponding ground-truth label. Here, $|X_{mn}|$ denotes the size of the set $X_{mn}$. The data, such as images of streets and pedestrians, can be generated by high-definition cameras inside and outside the vehicles. The steps involved in each iteration $r\in\left\{1,2,\ldots,R\right\}$ are as follows:
@@
  • Uploading and Aggregation: MUs transmit their local updates $\jmath_{mn}^{(r)}$ to the AI engine of MSP $n$. Upon receiving all the local updates, the MSP $n$’s AI engine aggregates them through the following weighted average:
@@
III-C Immersion Metric of Local Model
MSPs determine the rewards for MUs based on the immersion of the local model (IoM) metric. Correspondingly, MUs need to decide the IoMs of local models provided to MSPs by allocating their computational and communication resources to maximize the benefits. This process highlights the key role of “IoM” in the interactions between the two parties, serving as both an evaluation criterion and a basis for decision-making. Unlike traditional FL and incentive mechanisms, vehicular metaverse scenarios with immersion requirements necessitate consideration of both the contribution and freshness of the local model [^8] For example, an MSP may receive fresher local updates but with lower contributions, or less fresh local updates but with higher contributions. , which will collectively affect the immersive experience of AR services, i.e., whether virtual objects can be placed in the physical world accurately and promptly.
Driven by the above considerations, we design the IoM based on four key dimensions: the freshness and accuracy of the local model, as well as the amount and potential value of raw training data. The freshness is captured through age of information (AoI) $\Delta_{mn}$ [31], while the other three dimensions are jointly reflected in contribution prediction $I_{mn}$. To measure the value of a local model that MU $m$ brought to the vehicular metaverse service of MSP $n$, we define the IoM as
@@ III-C Immersion Metric of Local Model
III-C1 Contribution Prediction II
The contribution prediction $I_{mn}$ from MU $m$ to MSP $n$ is determined by the accuracy $\theta_{m}$ of the local model, the total amount of training data $\lfloor\frac{T}{\tau_{n}}\rfloor|X_{mn}|$, and the potential value $\omega_{mn}$ of local data. Prior work [32] characterized model contributions using a logarithmic function of the amount of local training data. However, evaluating the contribution solely by the training data size is one-sided, as there may be a large amount of redundant data or the validity of the local model cannot be guaranteed. It is more practical to comprehensively evaluate the amount of training data, model accuracy, and the potential value of local data; thus, we denote $I_{mn}$ as
@@ III-C1 Contribution Prediction II
where $\lfloor\frac{T}{\tau_{n}}\rfloor$ is the number of iterations that the task of MSP $n$ can be performed within $T$. $\epsilon$ and $\eta$ are system parameters from experiments as obtained in [33]. $\theta_{m}\in\left(0,1\right)$ characterizes the accuracy of local training as decided by the MU. Here, $\theta\rightarrow 0$ means that a high-accuracy local model is available, while $\theta\rightarrow 1$ implies that the accuracy of the local model is low. The local data potential value $\omega_{mn}$ describes the difference between model predictions and true labels, defined as
@@
where $\hat{z}_{i}$ denotes the inferred label from the current global model and $z_{i}$ is its ground-truth label. If the value of $\omega_{mn}$ is large, it indicates that the current global model is not working well on local data or there are unseen samples. Therefore, updating the model by training on these data may lead to better performance of the global model. If the value of $\omega_{mn}$ is small, the current data has little potential to improve the global model further since the knowledge of the data has already been learned or the data size is small. Note that to shorten the decision time, we use the potential value of the initial local data to be a proxy for the average potential value of all data within $T$. This is a reasonable approach when $T$ is not too long.
Figure 5: The AoI evolution of the local updates sent from MU $m$ to MSP $n$.
@@ III-C2 Age of Information $\Delta$
The whole FL process involves model training, uploading, and waiting, which together constitute a complete cycle of model updating. From the perspective of MSP $n$, the age evolution of the model contributed by MU $m$ is shown in Fig. 5. Here, the filled circles denote the time instances when MU $m$ starts training the local model. The intersections show the time instances when MSP $n$ receives the corresponding local updates, at which point the AoI drops to the lowest.
We denote that at time $t$, the generation time of the last update received by MSP $n$ from MU $m$ is $s(t)$. Then, the instantaneous AoI of MU $m$’s local model measured at MSP $n$, $\Delta_{mn}(t)$, is
@@
which shows the elapsed time since the generation of the latest local updates. A smaller instantaneous AoI means that the latest local updates received are fresher at that time.
To obtain a comprehensive understanding of the overall freshness of model updates over a certain period, we use average AoI to evaluate the freshness of the models. With the aid of Figure 5, the time-averaged AoI of the local model from MU $m$ to MSP $n$ can be expressed as
$\overline{\Delta}_{mn}$ $=\mathop{\lim}\limits_{Z\to\infty}\frac{1}{Z}\int_{0}^{Z}{{\Delta_{mn}}(t)}\,dt=\mathop{\lim}\limits_{Z\to\infty}\sum_{r=1}^{Z}\frac{{Q_{mn}}^{(r)}}{Z}$ (6)
@@
where ${Q_{mn}}^{(r)}$ is the $r$-th trapezium under the curve. A lower average AoI indicates that the local updates are generally fresher over a long period.
Since MU $m$ continuously collects data during the FL process, the amount of training data used in each round is $|X_{mn}|=x_{m}\tau_{n}$, where $x_{m}$ denotes the data (number of floats) collected per unit time. The cumulative time $T_{mn}^{c}$ for local training is determined by the amount of training data, computational resource $f_{mn}$ allocated to MSP $n$, and its local accuracy threshold $\theta_{m}$, which is equivalent to
@@
Furthermore, the upload time $T_{mn}^{t}$ depends on the size of local updates $b_{mn}$ (in bits) and the resource $B_{mn}$ (in Hz) allocated for communication between MU $m$ and MSP $n$ (i.e., uplink bandwidth), defined as
@@
where $\varsigma_{mn}=\frac{p^{t}_{m}g_{mn}}{\sigma^{2}}$ denotes the signal-to-noise ratio (SNR) [35] for the communication between MU $m$ and MSP $n$. $g_{mn}$ and $p^{t}_{m}$ are the corresponding channel gain and the transmission power of MU $m$, respectively. $\sigma^{2}$ is the power of additive white Gaussian noise (AWGN).
@@ IV-A Utility of MU
$\max\,\,\Phi_{m}(\bm{f}_{m},\bm{B}_{m})=\sum_{n=1}^{N}(p_{mn}V_{mn}-C_{mn})$ (11)
$\mathrm{s.t.}$ $\ C1:\sum_{n=1}^{N}f_{mn}<f_{m}^{max},\,\,\sum_{n=1}^{N}B_{mn}<B_{m}^{max},$
$\ C2:\log(1/\theta_{m})\frac{x_{m}\tau_{n}}{f_{mn}}+\frac{b_{mn}}{B_{mn}\log_{2}(1+\varsigma_{mn})}\leq\frac{1}{2}\tau_{n},$
@@
where $p_{mn}$ is the reward offered by MSP $n$ to MU $m$ for their contribution. $S_{m}$ indicates the minimum computational resource required to enjoy other basic services. Constraints C1 and C3 indicate that the MU’s total computational and communication resources are limited and cannot be fully used to contribute local models, given the MU’s demand for basic services. Constraint C2 ensures that $V_{mn}$ is non-negative, guaranteeing that the value derived from the task meets the MSP’s minimum requirements. Specifically, the total time for local training and uploading in each round cannot exceed the time constraint $\frac{1}{2}\tau_{n}$. The goal of each MU is to choose the optimal allocation of computational $\bm{f}_{m}$ and communication $\bm{B}_{m}$ resources to maximize its utility.
@@ IV-B Utility of MSP
Each MSP has a gain function $\psi$ associated with the IoM to quantify the cumulative benefit of all local models contributed by MUs. In this paper, we adopt $\psi_{n}=\mu_{n}\ln(1+\sum_{m=1}^{M}V_{mn})$ as a monotonically increasing, differentiable, strictly concave IoM function, which is a simplification of the widely adopted function [36]. Let $\bm{p}_{n}\triangleq\left(p_{1n},\ldots,p_{Mn}\right)$ and $\bm{p}_{\textrm{-}n}$ represent the reward profile of MSP $n$ and all other MSPs except MSP $n$, respectively. Then, the optimization problem for the MSP is defined as
@@
$\mathrm{s.t.}\ p_{mn}>0,n\in\mathcal{N},m\in\mathcal{M},$
where $\mu_{n}$ refers to the profit conversion coefficient [37] from IoM, which is adjusted according to the AR services offered by the different MSPs. Note that the other MSPs’ decisions $\bm{p}_{\textrm{-}n}$ are captured by the MUs’ responses. Moreover, when MSP $n$ determines its reward $\bm{p}_{n}$ for different MUs, the MSP needs to consider the rewards offered by other MSPs (i.e., $\bm{p}_{\textrm{-}n}$ ) as well as the strategies of all MUs (i.e., $(\bm{f}_{m},\bm{B}_{m}),m\in\mathcal{M}$). This thereby leads to reward competition among MSPs. Therefore, the optimization among MSPs can be considered a non-cooperative game, termed a multi-MSP rewarding game $\Omega$, as follows.